# Are language models rational? The case of coherence norms and belief revision

### v2

@@ 4 Belief revision norms for language models
The logical and probabilistic coherence norms we considered so far were synchronic norms: they apply to a system with beliefs at a time. Belief revision, on the other hand, is essentially diachronic and concerns how beliefs should be changed over time. The most natural extension of probabilism to a diachronic norm is Bayesianism, which we here take to be the view that credences should be updated in light of new evidence in accordance with Bayes’ Theorem. Rationality requires, according to a Bayesian, that one conditionalize on new evidence and adjust one’s credences accordingly over time. This is a natural rational norm for belief revision as it applies to us humans and similar creatures. But it is not clear how it applies to machine learning models like language models. First of all, it is unclear how to make sense of the notion of evidence for a language model. For us human beings, evidence is tied to perception by observing the world or to learning from others directly in communication. But language models by themselves do not have perceptual abilities. Even multi-modal models do not have perception as such. A model trained on language as well as images from comic books, say, would not thereby have a perceptual ability, since the images are not properly connected to the world. Such a model is multimodal, but that by itself does not make it perceptual. Perception-like abilities could in principle be integrated with language models, say via feeding it images from a camera that depicts the world, but without a clearer understanding of how that will go, it is hard to make sense of evidence for language models, and with it is hard to make sense of Bayesian updating and Bayesian approaches to belief revision. Understood as a diachronic norm of rationality, it does not apply to language models as they are at present.
@@
Language models nonetheless can change their beliefs: they can be fine-tuned to adopt a particular belief or whole set of beliefs, and they can be directly edited to change a particular belief. [^8] Zhu et al. (2020); De Cao et al. (2021); Dai et al. (2021); Hase et al. (2021) But such change of belief of the model is not a rational process for the model itself. It is simply an external modification of the model, and the question of rationality does not directly apply to it from the point of view of the model. It is a change that happens to it from an external source, not a direct response to rational forces by the model itself.
Nonetheless, the synchronic coherence norms that apply to the model will have an effect on what rational requirements there are for it after the change. Even if the change is not brought about by rational means, the resulting state after the change is subject to the synchronic norms of rationality like logical coherence and probabilism, assuming they apply to the model in the first place. The synchronic norms put rational pressure on the model’s present state. No matter how it got into that state, through rational means or merely external causal forces, the synchronic norms require it to be in a coherent state afterwards.
Thus for many present language models diachronic rational norms like Bayesian updating do not yet apply, even if the model is fine-tuned for truthfulness. However, synchronic norms do apply in this case, and they apply no matter how the model got into its present state: through model editing to correct a false belief, or fine-tuning, or through other means that change its beliefs. As long as the model continues to aim at truthfulness, synchronic coherence norms continue to apply. Future models might more directly be under diachronic rationality norms, but for that to be so something like a notion of evidence for the model must be made sense of.
@@ 5 Living up to the normative requirements
The question what norms apply to language models is one that can be discussed rather generally. One can argue, as we did above, that particular classes of models are under particular norms of rationality, while others are not. But the question whether a particular model lives up to the norms that apply to it is one that is specific to each model. It depends on how well the model is doing what it is supposed to do, and that depends on the details of the model, and can’t be determined in general. Nonetheless, there is a general question in the neighborhood, which is one about the effectiveness of rational requirements on models in general. Our human minds are under rational norms and somehow these norms are generally effective. Although we are often irrational, it is not the case that rational norms have no effect on us at all. They often push us in the right direction. How rational and irrational human beings are is a widely studied question. [^9] Tversky and Kahneman (1974) Our concern here is not human beings, but language models. And to determine how rational they are faces a special additional difficulty: Language models face a conflict of norms that does not seem to carry over to human rationality.
@@
Language models can be under competing normative requirements, and because of this it is hard to see how effective the norms of rationality are that apply to them. And for a similar reason, it is hard to see how rational they are when we try to assess the rationality of their beliefs. To illustrate the issue, consider a language model fine-tuned for truthfulness, but pre-trained on incoherent text. Such a model is under rational coherence norms, since it aims at the truth. In particular, it is under the norm to have coherent, logically consistent beliefs. To see if it lives up to this coherence norms we can empirically determine if it has consistent beliefs. We can query the model with questions and determine if it gives us logically consistent answers. And thus it would seem easy to test whether it lives up to the coherence norms that apply to it.
@@
But unfortunately it is not that easy. Suppose that we find that the model does not live up to this requirement and has inconsistent beliefs. Is the reason for this that it fails to live up to a normative requirement of logical coherence, which applies to it, and thus it is irrational? That is not so clear, since the model is not only fine-tuned for truthfulness, but also pre-trained on incoherent training data, i.e. on text that is not logically consistent. And as a machine learning model, it is also under the requirement to model the training data. These two requirements, however, are in conflict in this case. Being pre-trained on incoherent text might support an incoherent set of beliefs as the proper response. But being fine-tuned for truthfulness requires a coherent state as a response. If we encounter an inconsistent set of beliefs in a language model, is that because it fails to live up to the rational requirement of coherence, or is it because it does live up to the requirement of learning the training data? For language models like the ones we have today, it is hard to separate the two normative forces, and thus it is hard to empirically test whether rational norms are effective.
@@
To solve this problem, and to properly test how effective rational norms are in language models, we need to control one variable to test for the other. In particular, we need to control the incoherence of the training data in pre-training to test for rationality in the fine-tuned models. If we knew, say, that the model was only pre-trained on logically consistent training data, then any inconsistency in its beliefs is a result of a rational failure. And if we knew that if the training data was only slightly incoherent, then the degree of incoherence might correspond to a degree of irrationality. But, of course, such experiments are not possible with present language models, since they are trained on a large corpus of text that is not controlled for coherence, and in fact is known to be vastly incoherent. And what is worse, it is known to be hard to test any large body of natural language text for logical coherence. Just to test for Boolean satisfiability is an NP-complete problem (Cook, 1971). It is likely that pre-training work involving synthetic corpora will be required to determine whether language models produce inconsistencies as a result of adherence to their training data or a failure of rationality.
In related empirical work, Hase et al. (2024), we hope to make progress on these questions in just this way: by testing rationality in a new language model that is pre-trained from scratch on a novel semi-synthetic dataset controlled for coherence. In particular, we hope to assess probabilistic coherence in models pre-trained on datasets that are incoherent in a measurable way. To do this properly, we need to work with semi-synthetic data whose coherence is controllable in a precise way. In a nutshell, it appears that the model’s internal probabilities track the degree of incoherence in the training data, which largely speaks in favor of rational norms being effective. However, when it comes to model updating and belief revision, present methods fall far short of an ideal Bayesian agent. Again, a full discussion of this empirical work is in Hase et al. (2024).
@@ 6 Conclusion
To determine whether language models are rational, irrational, or arational, is a crucial part in understanding them, both internally and why they produce the outputs they produce. In this paper we argued that this question does not have a uniform answer for all language models. In particular, we argued that pre-trained models generally are arational, since their internal states are solely under the norms of machine learning models to learn the training data. However, fine-tuned models can be under rational norms, as long as they are fine-tuned for truthfulness, which is the case for models fine-tuned with RLHF, or on special datasets selected for truthfulness. Such models are under both logical and probabilistic coherence norms. With regards to the latter, we proposed that the notion of a credence makes sense for language models, and that we can define a coherent credence function for a model from next token probabilities, exploiting the minimal assent connection MAC. As for how well language models live up to coherence norms, we leave this question to future work studying model coherence with the proper controls for the consistency of the training data.
Overall, this work points in the direction that considerations of rationality do apply to language models, although the difference of being trained on incoherent data in pre-training, and not a coherent world directly, makes an important contribution to understanding how rational norms have a different pull in language models than in humans. And even if it is not clear how well language models live up to coherence norms in practice, we can at least employ our tools of predicting and explaining behavior in terms of representational states for understanding both humans and models.

### v3

@@ 4 Belief revision norms for language models
The logical and probabilistic coherence norms we considered so far were synchronic norms: they apply to a system with beliefs at a time. Belief revision, on the other hand, is essentially diachronic and concerns how beliefs should be changed over time. There are different approaches to characterizing the norms that govern such rational change of belief. One is focused on full belief and a logical system of update rules that aim to capture such change. [^12] See alchourron1985logic for a paradigmatic example. However, logic-based approaches are not naturally integrated with language models, as they do not operate primarily through symbol manipulation. It is much more natural when considering belief revision in language models to focus on norms that operate on credences, since, as we have seen above, those can be directly associated with belief in language models. This is what we did in the last section with synchronic coherence norms like probabilism. The most natural extension of probabilism to a diachronic norm is Bayesianism, which we here take to be the view that credences should be updated in light of new evidence in accordance with Bayes’ Theorem. Rationality requires, according to a Bayesian, that one conditionalize on new evidence and adjust one’s credences accordingly over time.
This is a natural rational norm for belief revision as it applies to us humans and similar creatures. But it is not clear how it applies to machine learning models like language models. First and foremost, it is unclear how to make sense of the notion of evidence for a language model. For us human beings, evidence is naturally tied to perception by observing the world or to learning from others directly in communication. But language models by themselves do not have perceptual abilities. Even multimodal models do not have perception as such. A model trained on text as well as images from comic books, say, would not thereby have a perceptual ability, since the images are not properly connected to the world. Such a model is multimodal, but that by itself does not make it perceptual. Perception-like abilities could in principle be integrated with language models, say via feeding it images from a camera that depicts the world, but without a clearer understanding of how that will go, it is hard to make sense of evidence for language models and with it is hard to make sense of Bayesian updating and Bayesian approaches to belief revision. Understood as a diachronic norm of rationality, it does not seem to apply to language models as they are at present.
@@
Language models nonetheless can change their beliefs: they can be fine-tuned to adopt a particular belief or whole set of beliefs, and they can be directly edited to change a particular belief.131313zhu2020modifying; de2021editing; dai2021knowledge; hase2021language But such change of belief of the model is not a rational process for the model itself. It is simply an external modification of the model, and the question whether that change in belief was rational does not directly apply from the point of view of the model. It is a change that happens to it from an external source, not a direct response to rational forces like evidence. But the question remains how the model should update its beliefs and credences in other propositions given that a particular belief has been changed. If a belief changes, then this will likely lead to incoherence and inconsistency unless other beliefs change as well. To preserve coherence, any change in belief cannot just be local and isolated, it must also affect many other beliefs that are connected to that belief. Such connections can come in the form of logical relationships or probabilistic ones, for example, conditional credences.
Thus, even when a belief change happens in a non-rational way, and not as the result of responding to evidence, there are other norms of rationality that require further subsequent belief changes. Those norms of rationality can require that, given a particular new belief, various other beliefs need to be adjusted in light of how they are connected to the new belief. If the model now believes that the butler did it, then it should stop believing that the maid did it. And if it increases its credence that the cook did it, then it should decrease its credence that one of the other people did it.
This highlights that synchronic coherence norms generate rational requirements tied to belief change, even if that belief change was not itself done rationally. Even if the model is externally modified to hold a certain new belief, by fine-tuning or direct editing, the synchronic coherence norms require further changes in belief and credence to restore coherence or achieve it outright. Synchronic coherence norms can thus support conditional diachronic coherence norms, even if the model does not have the ability to respond to evidence. Conditional on having a given new belief, synchronic coherence norms require various revisions to its other old beliefs as well.
What then do these norms demand for how the model’s beliefs are to be modified? In general, there is no unique answer to this question. The easiest way to achieve coherence again after a belief change, assuming that coherence was there in the first place, is simply to undo the belief change and revert back to one’s original state. But suppose that we keep the externally induced belief change and we will not revert back. Given that we fix the new belief and its credence, what do synchronic coherence norms require of the model to change as a consequence of this new fixed belief? What else needs to change, and how, to live up to the synchronic coherence norms?
The most natural way to achieve coherence while keeping the new belief B fixed is to also keep conditional credences fixed as much as possible and adjust one’s other beliefs in light of these conditional credences and the new fixed belief B. How this change should happen in a broadly Bayesian framework will depend on what credences one has in the new belief B. If it is considered fixed and certain, then B functions in effect like Bayesian evidence and one simply needs to adjust one’s credences in any other belief A in accordance to one’s prior conditional credences in A given B. That is to say:
$P_{new}(A)=P_{old}(A|B)$ (6)
If we, on the other hand, treat B not as certain and give it a regular credence below 1, then the most natural way to achieve coherence is to weigh the conditional credences relating A and B by the new credence in B. This corresponds to Jeffrey conditionalizing:
$P_{new}(A)=P_{old}(A|B)P_{new}(B)+P_{old}(A|\neg B)P_{new}(\neg B)$ (7)
These are widely discussed methods for updating one’s credences teitelbaum:bayes, and both are, of course, not without their critics. We do not have to settle the issue here as to which method is best. What matters for us here is not which method precisely is the correct one, but rather that there are derivative broadly Bayesian diachronic norms connected to belief revision that apply to language models and which are derivative on the synchronic norms defended above. Even though language models of today cannot yet be said to respond to evidence, nonetheless belief revision norms make sense for them. Externally induced belief changes via fine-tuning or other model editing methods bring with them diachronic norms that govern how the model should adjust its beliefs and its credences. The details of these norms can be debated, but Jeffrey conditionalization as in (7) is a very plausible candidate.
Having thus defended both synchronic as well as diachronic norms for belief in language models, it is now time to address the third and final of our initial questions: do present day language models live up to the norms of rationality that apply to them?
@@ 5 Living up to the normative requirements
The question what norms apply to language models is one that can be discussed rather generally and abstractly. One can argue, as we did above, that particular classes of models are under particular norms of rationality, while others are not. But the question whether a particular model lives up to the norms that apply to it is one that is specific to each model. It depends on how well the model is doing what it is supposed to do, and that depends on the details of the model, and cannot be determined in general. Nonetheless, there is a general question in the neighborhood, which is one about the effectiveness of rational requirements on models in general. Our human minds are under rational norms and somehow these norms are generally effective. Although we are often irrational, it is not the case that rational norms have no effect on us at all. They often push us in the right direction. How rational and irrational human beings are is a widely studied question (tversky-kan1974). Our concern here is not human beings, but language models. And to determine how rational they are faces a special additional difficulty: Language models face a conflict of norms that does not seem to carry over to human rationality.
@@
Language models can be under competing normative requirements, and because of this it is hard to see how effective the norms of rationality which apply to them are. For a similar reason, it is hard to see how rational they are when we try to assess the rationality of their beliefs. To illustrate the issue, consider a language model fine-tuned for truthfulness but pretrained on incoherent text. Such a model is under rational coherence norms since it aims at the truth. In particular, it is under the norm to have logically consistent beliefs. To see if it lives up to this coherence norm we can empirically determine if it has consistent beliefs. We can query the model with questions and determine if it gives us logically consistent answers. And thus, it would seem easy to test whether it lives up to the coherence norms that apply to it.
@@
But unfortunately, it is not that easy. Suppose that we find that the model does not live up to this requirement and has inconsistent beliefs. Is the reason for this that it fails to live up to a normative requirement of logical coherence, which applies to it, and thus it is irrational? That is not so clear, since the model is not only fine-tuned for truthfulness, but also pretrained on incoherent training data, i.e. on text that is not logically consistent. And as a machine learning model, it is also under the requirement to model the training data. However, these two requirements, logical coherence and learning the training data, are in conflict in this case. Being pretrained on incoherent text might support an incoherent set of beliefs as the proper response. But being fine-tuned for truthfulness requires coherent beliefs. If we encounter an inconsistent set of beliefs in a language model, is that because it fails to live up to the rational requirement of coherence or is it because it does live up to the requirement of learning the training data? For language models like the ones we have today, it is hard to separate the two normative forces and settle which one should have stronger pull. Thus, it is hard to empirically test whether rational norms are effective.
@@
To solve this problem, and to properly test how effective rational norms are in language models, we need to control one variable to test for the other. In particular, we need to control the incoherence of the training data in pretraining to test for rationality in the fine-tuned models. If we knew, say, that the model was only pretrained on logically consistent training data, then any inconsistency in its beliefs is a result of a rational failure. And if we knew that the training data was only slightly incoherent, then the degree of incoherence in the data might correspond to a degree of irrationality in the model. But, of course, such experiments are not possible with present language models, since they are trained on a large corpus of text that is not controlled for coherence and in fact is known to be vastly incoherent. And what is worse, it is known to be hard to test any large body of natural language text for logical coherence. Just to test for Boolean satisfiability is an NP-complete problem (cook-complexity). It is likely that pretraining work involving synthetic corpora will be required to determine whether language models produce inconsistencies as a result of adherence to their training data or a failure of rationality.
We take up this question in related empirical work (hase2024fundamentalproblemsmodelediting). The goal of the experiments there is two-fold. We test (1) whether language models live up to synchronic coherence norms, with respect to a synthetically constructed pretraining corpus, and (2) whether they live up to derivative diachronic norms (see Sec. 4), including the fulfillment of synchronic norms before and after editing a language model to adopt a new fact about the world.
To test for (1), we pretrain a language model on 1 billion tokens of synthetic data constructed algorithmically according to the rules of a formal language. The corpus consists of sentences about a hypothetical world. There are a limited number of true, atomic propositions about this world. We construct a partially incoherent dataset by including true sentences as well as false sentences (sentences that contradict the true sentences) at known frequencies. Crucially, this enables us to obtain a “ground truth” credence for each atomic sentence in the data. In fact, this “ground truth” credence is given to us by a Bayesian model, also fit to the incoherent corpus. We will refer to the probabilities learned by this Bayesian model as gold standard credences. Thus, to measure synchronic coherence, we compare the credences of our pretrained language model against these gold standard credences. When we pretrain a language model on our corpus, we find that the resulting language model is quite successful at generating the most common sentences, which are also the “true” sentences about our hypothetical world. But the model is quite unsuccessful at giving calibrated probabilities for these sentences. This means that the model’s credences are directionally correct, i.e. they track the truth, but generally imprecise, i.e. they diverge from the ideal Bayesian gold standard.
To test for (2), we take this pretrained model and fine-tune it to adopt a new fact about the world, represented by a sentence that contradicts what was seen in the pretraining data. Does this new model maintain a coherent picture of the world? We find that, no, these updates to the model immediately produce some incoherent beliefs. We can say so by assessing the updated language model against gold-standard posterior credences given by our Bayesian model, also fed the same new sentence. The Bayesian model performs exactly the kind of Jeffrey conditionalization given in Eq. 7, with the new sentence serving the role of evidence $B$. To give an example of the language model’s incoherence, when the model successfully learns the new sentence P, it does not then assign the appropriate credence to the sentence “not P”. It also does not assign the right credence to a sentence that is implied by P (along the lines of “X is a bachelor” implying “X is unmarried”). More details are available in our paper (hase2024fundamentalproblemsmodelediting).
In sum, this suggests that a pretrained language model’s credences only partially track the degree of incoherence in the training data. A model can learn what atomic propositions are true or not, but its credences obtain high Brier scores (i.e., high error). Moreover, coherence norms may sometimes have no effect. After fine-tuning on new sentences, these norms fall quite short of being effective.
These experiments are conducted in a controlled setting with a synthetic corpus in order to tease apart the role of pretraining data coherence and rational norms. To add context to this result, we note that larger-scale models (LLMs) trained on natural language text often appear to demonstrate a robust understanding of the world. We point to (wilie2024belief) for a study of how such LLMs perform logical inference and belief revision in a more naturalistic setting. We note that LLM abilities show limitations under these quantitative evaluations but have been growing more capable over time.
@@ 6 Conclusion
To determine whether language models are rational, irrational, or arational, is a crucial part in understanding them, both internally and why they produce the outputs they produce. In this paper we argued that this question does not have a uniform answer for all language models. In particular, we argued that pretrained models generally are arational, since their internal states are solely under the norms of machine learning models to learn the training data. However, fine-tuned models can be under rational norms, as long as they are fine-tuned for truthfulness, which is the case for models fine-tuned with RLHF, or on special datasets selected for truthfulness. Such models are under both logical and probabilistic coherence norms. With regard to the latter, we proposed that the notion of a credence makes sense for language models, and that we can define a coherent credence function for a model from next token probabilities, exploiting the minimal assent connection MAC. As for how well language models live up to coherence norms, we leave this question to future work studying model coherence with the proper controls for the consistency of the training data.
Overall, this work points in the direction that considerations of rationality do apply to language models, although the difference of being trained on incoherent data in pretraining, and not a coherent world directly, makes an important contribution to understanding how rational norms have a different pull in language models than in humans. And even if it is not clear how well language models live up to coherence norms in practice, we can at least employ our tools of predicting and explaining behavior in terms of representational states for understanding both humans and models.